OpenStudy (anonymous):

Find the difference. 7/10 - 3/10

OpenStudy (anonymous):

Is it 1/2?

OpenStudy (anonymous):

4/10

OpenStudy (anonymous):

its 2/5

OpenStudy (anonymous):

see how I got that?

OpenStudy (anonymous):

No :S

OpenStudy (anonymous):

7/10 -3/10 = 4/ 10 since the denominators are the same you just subtract the numerators which gets 7-3=4 put that over 10 and reduces to 2/5

OpenStudy (phi):

Here is a short video on subtracting fractions http://www.khanacademy.org/math/arithmetic/fractions/v/subtracting--fractions

OpenStudy (anonymous):

thats what school is for man